OpenStudy (anonymous):

using standard equation to give the radius and center of the circle; (x-4)^2+(y-3)^2=16 a little HELP!! plz!!

OpenStudy (amistre64):

that 16 tells us the radius is sqrt(16) = 4 the center of the circle is shown by whatever those numbers are that are messing around with your x and y parts. center is at(4,3)

OpenStudy (amistre64):

(x-4) says that it was originally at another location but we moved it to get a normal; reading.... we moved it by -4 to get it back to 0, so it was originally at 0+4 = 4

OpenStudy (anonymous):

Well you know the standard formula is (x-h)^2 + (y-k)^2 = r^2. In this case, r^2=16 so r=4. H and K are the x- and y- coordinates for the circle respectively, so you know from your equation that the x-coordinates of the center is 4 and the y-coordinate of the center is 3. So your center would be located at (4,3)

OpenStudy (anonymous):

can you please show me the steps PLEASE!!

OpenStudy (amistre64):

say you wanna look at something that is on a shelf, what do you do? you pick it up, move it over to where you can get a good lok at it, and when your done you put it back right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

well, when we look at circles, or any equation, we want to observe them from (x=0,y=0) so if they are over someplace else we gotta move them...

OpenStudy (amistre64):

lets say this circle is centered at (4,0) how far do we move it to get it to (0,0) so that we can look at it?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

no -3

OpenStudy (amistre64):

thats correct, so we make a note to ourselves going, we moved x -4, so that when we want to put it back....we can remember where we got it from. -4 was correct

OpenStudy (amistre64):

you see the (x-4) part in your problem?

OpenStudy (anonymous):

so is that how it came??

OpenStudy (amistre64):

thats exactly how it came; ... :)

OpenStudy (anonymous):

so if the equation is (x-5)^2+(y-1)^2=25 then r=5 and centre= (5,1)???

OpenStudy (amistre64):

you got it, that is absolutely correct..... good job :)

OpenStudy (anonymous):

so there is no complicated algebra we gotta do??

OpenStudy (anonymous):

so what if the equation is (x-(1/2))^2+(y+(3/4)^2=(1/4)

OpenStudy (amistre64):

then just pull out the fraction stuff as your answer, x = 1/2 y = -3/4 radius = sqrt(1/4) = 1/2

OpenStudy (anonymous):

ok so now i have a graph and i have to give the coordinates of the centre, the radius, and the equation of the circle the graph is like this:: . . . 5 . . . . . . . 4 . . . . . . . 3 . . . . . . . 2 . . . . . . . 1 . . . . . . . 0 . . . . . . -1 . . . . . . -2 . . . .